Prove by induction that
1+2+3+.....+n = n(n+1)÷2
1
Expert's answer
2021-09-22T16:28:50-0400

1) base case: n=1

for n=1, the statement take the next form: "1=1(1+1)\/2" . After simplification we get "1=1", which means the base case is true.

2) Induction step: under the assupmtion that statement is true for n=k, we will prove that it is true for n=k+1.

for n=k we get "1+2+3+...+k=k(k+1)\/2"

for n=k+1 we get:"1+2+3+...+k+(k+1)= k(k+1)\/2 + k+1= (k(k+1) + 2(k+1))\/2 = (k+1)(k+2)\/2"

The statement has been proven.